OpenStudy (anonymous):

Can someone please help me with this project, I could get an f in math and fail my grade if it isnt finished before my teacher grades papers

OpenStudy (anonymous):

There are only 3 questions, but I want to understand how to solve them too

OpenStudy (anonymous):

bruh,whats your project

OpenStudy (anonymous):

ok

OpenStudy (godlovesme):

can u post the questions :) i'll try @Yazmyn

OpenStudy (anonymous):

file:///C:/Users/daughter/Downloads/8MA_Edmodo_November2014.doc.pdf

OpenStudy (anonymous):

Sorry lmao

OpenStudy (anonymous):

They're on the link luv @Godlovesme

OpenStudy (anonymous):

use the @ sighn

OpenStudy (anonymous):

@openstudyhelp98765432 what kind of a helping tool are you?

OpenStudy (anonymous):

What kind of an answer is this? This would be the kind of answer Barney would use.

OpenStudy (godlovesme):

@Yazmyn the link won't open :(

OpenStudy (anonymous):

i ment use @ so i could go to it smart mouth

OpenStudy (anonymous):

http://prntscr.com/5rqfpf

OpenStudy (anonymous):

@Godlovesme @openstudyhelp98765432

OpenStudy (godlovesme):

@yazmyn r u there?

OpenStudy (anonymous):

im here

OpenStudy (godlovesme):

k so u r supposed to write the equation in slope-intercept form?

OpenStudy (anonymous):

I dont think it matters, it just has to be a problem that represents the cost of strawberries

OpenStudy (godlovesme):

okay so do u c any points that meets perfectly like in the strawberry line we have (1,2)

OpenStudy (anonymous):

I see (6,4)

OpenStudy (anonymous):

(6,3)

OpenStudy (godlovesme):

(x,y) that's how u write is so u mean (3,6) ?

OpenStudy (anonymous):

yea lol

OpenStudy (godlovesme):

:) how many up and over do u have to go to get another point?

OpenStudy (anonymous):

Up 2 over 1

OpenStudy (anonymous):

Help me with mine too @Godlovesme + @Yazmyn Solve. Remember to write it as an ordered pair!! Will fan and medal. (I won't forget!!) x+y=20. 10x +12y =224

OpenStudy (anonymous):

is it a set?

OpenStudy (anonymous):

@Multimedia8989

OpenStudy (anonymous):

Oh nvm

OpenStudy (anonymous):

Yazymn please help me okay? I'm begging you!

OpenStudy (anonymous):

Options?

OpenStudy (godlovesme):

cool slope=rise/run = 2/1 so the eqn will be y=2x confused @yazmyn @Multimedia8989 I'll try after i finsh this one :)

OpenStudy (anonymous):

If it says x=4 and y=45, just substitue the number in for the variables

OpenStudy (anonymous):

No im ok :)

OpenStudy (godlovesme):

can u try for the blueberries :P

OpenStudy (anonymous):

k

OpenStudy (anonymous):

up 3 oiver 2

OpenStudy (godlovesme):

yes what about the eqn rise/run =>

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

Idk what you did in the last one to get y=2x, would this one be y=3x?

OpenStudy (godlovesme):

sorry guys my internet was down but the answer should be y=3/2x

OpenStudy (godlovesme):

@yamyn & @Multimedia8989 if u need help open a new question and tag me :)

OpenStudy (anonymous):

Ooohh

OpenStudy (anonymous):

So Strawberries= y=2x and blueberries= y=3/2x @Godlovesme

OpenStudy (godlovesme):

Yes :D

OpenStudy (anonymous):

Ok so now we put each in an equation?

OpenStudy (godlovesme):

Yes u already got the eqns right?

OpenStudy (anonymous):

is y=3/2x an eqsn?

OpenStudy (godlovesme):

Yep for the blueberries

OpenStudy (anonymous):

ooohh yaay first part done

OpenStudy (anonymous):

Now it's which fruit is cheaper and how can you tell?

OpenStudy (godlovesme):

Then can u compare the cost

OpenStudy (godlovesme):

Well one costs$2 and the other costs$1.50 sooo...

OpenStudy (godlovesme):

@yzmyn hope that helped And thanks for the medal :)

OpenStudy (anonymous):

Which one is 1.50?

OpenStudy (anonymous):

Youre welcome :)

OpenStudy (godlovesme):

remember 3/2 for the blueberries :P @Yazmyn

OpenStudy (anonymous):

So tht makes the strawberry $2

OpenStudy (godlovesme):

yep which one is cheaper then?

OpenStudy (anonymous):

Blu lmao

OpenStudy (anonymous):

Can u help me with the last part tomorow?

OpenStudy (godlovesme):

idk can i help u now :)

OpenStudy (anonymous):

yea lol

OpenStudy (godlovesme):

sup? :P

OpenStudy (anonymous):

1 sec

OpenStudy (godlovesme):

k